# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

### v1

@@ 4.1 Threat Model
Following previous work [17], we define a threat model to describe the goals and capabilities of adversaries and T2I service providers, thereby clarifying the scope of our proposed defense framework.
Goals and Capabilities of Adversaries. The primary goal of adversaries is to bypass the safeguards in T2I models and induce the models to generate high-quality inappropriate or restricted content. We assume that adversaries have the capability to locally deploy an open-source T2I model, although this model fails to meet quality requirements. Adversaries can utilize the local model to create or collect any adversarial prompts, such as searching for adversarial text through optimization techniques. Additionally, we assume that adversaries can interact with the target model by attempting various prompt variations and can understand the security measures equipped within the target model, but they cannot directly access the internal parameter information of the target model.
Goals and Capabilities of T2I Service Providers. The primary goal of T2I service providers is to prevent unsafe generation and ensure the quality of the generated images. In terms of capabilities, T2I service providers have complete access to the parameters of the T2I model, enabling them to optimize additional modules or update model parameters based on gradient information. Furthermore, service providers can integrate various safeguards, such as Negative Prompt, aimed at enhancing the safety of the T2I model from a visual perspective.
@@ 4.2 Overview
Based on the analysis of the properties of safety generation methods, our goal is to develop a safe generation framework that embodies these properties, providing comprehensive defense for T2I models in complex real-world scenarios to ensure generated content adheres to usage policies.
Key Idea. A T2I model essentially functions as a text-driven image generator, with its primary objective being to generate high-quality images that align with the given prompt. Ensuring semantic consistency between the input text and the output image is a core requirement of the T2I model. For a well-trained, normal T2I model, it is challenging to generate an unsafe image from a clean prompt that contains no inappropriate descriptions. For instance, it is nearly impossible for a T2I model to produce an image containing nudity when prompted with ”A photo of a cute dog”, as this would clearly violate consistency requirements. In other words, any inappropriate content in the generated image likely originates from inappropriate descriptions within the input prompt, suggesting that a T2I model’s primary security risks may arise from the text encoder rather than other components.
Based on this insight, we introduce Embedding Sanitizer (ES), a vision-agnostic safe generation defense framework designed as an add-on plugin that operates on prompt embeddings, aiming to erase target concepts at the source (i.e., prompt embeddings). This approach to concept erasure provides ES with two practical advantages. First, most existing safe generation frameworks focus on removing target concepts from visual representations—typically by modifying the denoiser or denoising process—while ignoring the importance to remove toxic representations within text embeddings. This characteristic renders ES independent of other safe generation frameworks, providing it with compatibility and allowing seamless integration with various safeguards. Second, ES ’s training process only requires access to the text encoder, without needing access to the denoiser or image encoder. This design enables ES to function as a plug-and-play module that can be effortlessly transferred across different T2I models.
Next, we formalize the key idea of ES. Assume that the T2I model to be fortified is equipped with a text encoder denoted as $\mathcal{F}_{t}$. Given a toxic prompt $P_{t}$ containing inappropriate descriptions, such as A nude person, ES aims to transform this into a clean prompt $P_{c}$ with similar semantics, such as A dressed person. This can be represented as:
$\text{ES}(\mathcal{F}_{t}(P_{t}))\approx\mathcal{F}_{t}(P_{c}),$ (2)
where ES takes the toxic prompt embedding $\mathcal{F}_{t}(P_{t})$ as input and performs sanitization to bring the output embedding close to that of the clean prompt $P_{c}$.
@@
Overall Pipeline. The entire execution process of ES is illustrated in  Figure 2. At a high level, when a user inputs a prompt to query the T2I (text-to-image) model, the text encoder first uses a tokenizer and embedding lookup to encode each token in the text prompt into embeddings. The text encoder then further processes these embeddings through a self-attention mechanism to extract high-level feature representations, which are fed as input to ES. Finally, ES performs a sanitization process on the embeddings, aiming to erase inappropriate concepts and transform them into clean embeddings devoid of any unsafe elements.
At a microscopic level, the embeddings output by the text encoder are first processed by EmbeddingNet (E-Net), which is designed to capture toxic embeddings containing inappropriate concepts within the original suspicious embeddings. ScoreNet (S-Net) then takes the toxic embeddings as input, assigning a score to each token constrained within the range [0, 1] by a Sigmoid activation function. For toxic tokens, S-Net is inclined to predict a higher score to strongly erase the target concept. For clean tokens, S-Net tends to assign a lower score, aiming to minimize interference with the generation of non-target concepts. The scores predicted by S-Net are multiplied with the toxic embeddings predicted by E-Net, and the suspicious concepts undergo sanitization by subtracting the toxic components from the suspicious embeddings in a residual connection. The sanitized toxic embeddings can then serve as an approximate representation of clean embeddings, providing guidance for generating appropriate content.
It is worth noting that ES only sanitizes embeddings for positive prompts and does not modify embeddings for negative prompts. This preserves the text encoder’s ability to represent inappropriate concepts, thereby not interfering with the operation of negative prompt-based safe generation frameworks, such as SLD.
4.3 Architectural Design
Designed with modularity in mind, ES functions as an independent component and can seamlessly integrate with other safeguards. This modular approach enables ES to serve as a plug-and-play module compatible with various T2I systems that share a same text encoder. As shown in Figure 2, ES’s architecture comprises two main components: E-Net and S-Net, which collaboratively perform the crucial sanitization function. Below, we provide a detailed description of each component’s architectural design and explain how these designs confer interpretability and controllability to ES. The model hyperparameter configurations for these two components are summarized in Table II.
@@ 4.3 Architectural Design
E-Net. The design of E-Net is based on the Transformer encoder architecture and includes six attention modules to efficiently capture potential inappropriate content in the input prompt. Each attention module employs an 8-head multi-head attention mechanism to enable E-Net to semantically analyze various parts of the input, allowing for a deeper understanding of tokens containing inappropriate concepts. E-Net’s role is to capture harmful concepts from the original suspicious embedding, producing a set of toxic embeddings that represent inappropriate content. Suppose $P_{s}$ is a suspicious prompt potentially containing inappropriate concepts, and it is encoded by the text encoder $\mathcal{F}_{t}$ to obtain the initial suspicious embedding, denoted as $\text{Emb}_{s}=\mathcal{F}_{t}(P_{s})$. E-Net, represented by $\mathcal{F}_{\theta}$, takes these embeddings as input and extracts identifiable toxic embeddings $\text{Emb}_{t}$, which can be formulated as:
@@
S-Net. S-Net is a multilayer perceptron (MLP) consisting of two fully connected layers, with its input dimension matching the output dimension of the embedding produced by E-Net. To constrain its output to [0, 1], a Sigmoid activation function is applied. The primary function of S-Net is to assign a score to each token in the input prompt, evaluating its potential harmfulness and indicating its contribution to the generation of inappropriate content, thereby providing ES with interpretability. Ideally, S-Net assigns high scores to toxic tokens, such as nude, enabling stronger erasure of toxic embeddings during the subsequent sanitization step. On the other hand, it assigns low scores to benign tokens, such as dog, minimizing disruption to the generation of non-toxic content. Represented by $\mathcal{F}_{\phi}$, S-Net takes the toxic embedding extracted by E-Net, $\text{Emb}_{t}$, as input and predicts a set of scores $\mathcal{S}$ for each token, expressed as:
@@
Embedding Sanitization. In the final sanitization step, the toxic embedding predicted by E-Net are multiplied by the scores predicted by S-Net, dynamically adjusting the sanitization strength according to each token’s score. This adjustment strategy provides two main advantages for ES ’s performance. First, it ensures that tokens identified as more toxic (i.e., with higher scores) undergo stronger sanitization, thereby enhancing the erasure of inappropriate content. Second, it preserves appropriate content by applying weaker sanitization to tokens identified as non-toxic (i.e., with lower scores). Next, ES sanitizes the original suspicious concept by subtracting (erasing) the toxic components from the suspicious embedding $\text{Emb}_{s}$ via residual connection to obtain a clean embedding $\text{Emb}_{c}$ that no longer contains inappropriate concepts. The sanitization process is formulated as:
$\text{Emb}_{c}$ $=\mathcal{F}_{t}(P_{s})-\alpha\cdot\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))$ (5)
@@
It is worth noting that we introduce a hyperparameter $\alpha$ to control the strength of sanitization in the inference phase. When $\alpha$ approaches 0, sanitization is weak, and ES ’s effect is minimal, resulting in $\text{Emb}_{c}$ being close to the output of the text encoder $\text{Emb}_{s}$. Conversely, as $\alpha$ increases, sanitization strength is significantly enhanced, effectively erasing harmful components from $\text{Emb}_{s}$ and producing a cleaner embedding $\text{Emb}_{c}$. A smaller $\alpha$ reduces interference with the input prompt, preserving more of the original semantics but with weaker sanitization, while a larger $\alpha$ enhances sanitization at the cost of some semantic preservation.
Therefore, $\alpha$ provides ES with controllability, enabling it to adjust sanitization strength flexibly according to the safety requirements of specific application scenarios, balancing between retaining content relevance and effectively erasing inappropriate concepts.
@@
4.4 Training Details
ES is essentially an end-to-end deep learning model that takes the original prompt embedding as input and outputs the sanitized clean embedding, which in turn guides the T2I model to generate appropriate content. Below, we introduce the training process of ES in detail, including training data preparation and objective function customization.
@@ 4.4 Training Details
Training Data Preparation. To sanitize prompt embeddings, we first select an anchor concept corresponding to the target concept to be erased, providing ES with a clear direction for sanitization. This approach enables us to frame the embedding sanitization as a transfer problem toward the anchor concept. When the T2I model receives a prompt containing a target concept, ES redirects this concept toward the predefined anchor concept to achieve embedding sanitization. We design three approaches to construct $\langle$ target concept, anchor concept $\rangle$ pairs:
@@
  • Antonyms of the target concept, e.g., $\langle$ illegal, legal $\rangle$;
@@
  • Hypernyms of the target concept, e.g., $\langle$ blood, liquid $\rangle$;
  • A neutral placeholder, e.g., $\langle$ weapon, __\__ $\rangle$.
@@
With the default settings we tend to use antonyms as anchor concepts to provide a more effective sanitization direction for the target concept.
@@
Considering that the performance of deep learning models largely depends on vast amounts of training data, we design an automated method for synthesizing prompts to provide ES with an almost limitless supply of training samples. Specifically, we view the synthetic prompt $P$ as a sequence of $n$ tokens drawn from the vocabulary $\mathcal{V}$, represented as $P=\{\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n}\}$. During the data synthesis process, we first convert the target concept set $\mathcal{C}_{t}=\{c_{t}^{1},c_{t}^{2},\ldots,c_{t}^{m}\}$ containing $m$ concepts and the anchor concept set $\mathcal{C}_{a}=\{c_{a}^{1},c_{a}^{2},\ldots,c_{a}^{m}\}$ into their corresponding token IDs using a tokenizer. Next, within the maximum input length allowed by the text encoder (e.g., SD-v1.4 allows a maximum of 77 tokens), we randomly select a number of tokens to compose $P$, expressed as $P=\{(\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n})\mid\text{token}_{i}\in\mathcal{V},\forall i=1,2,\ldots,n,n\geq 3\}$. Here, the requirement that $n\geq 3$ ensures that $P$ contains at least three tokens: a starting token, a random token, and an ending token, which guarantees that $P$ complies with input specifications and is not empty. Finally, we select random positions excluding the first and last tokens to replace them with random concepts from $\mathcal{C}_{t}$, thereby forming a toxic prompt containing a random number of target concepts, denoted as $P_{t}$. We then replace the target concepts in $P_{t}$ with their corresponding anchor concepts to create the clean prompt $P_{c}$. This dataset construction strategy enables ES to learn how to transform harmful prompts $P_{t}$ into clean prompts $P_{c}$, ensuring that the generated images reflect appropriate content without losing contextual relevance.
The example in Figure 3 provides an intuitive illustration of a randomly synthesized $P_{t}$ and its corresponding $P_{c}$, where red tokens indicate target concepts and green tokens represent anchor concepts. Automated data generation strategy significantly enhances the generalization capability and adaptability of ES. First, synthetic data allows us to generate a vast number of prompts containing target concepts in various contexts, assisting ES in learning to identify and erase inappropriate content across diverse scenarios. Second, the generation of synthetic data enables precise control over the combinations of target concepts and anchor concepts, allowing ES to effectively adapt to different application requirements. Furthermore, generating training data in an automated manner can substantially reduce manual annotation costs and mitigate dependence on large-scale real-world inappropriate datasets, ensuring that ES can be trained broadly while safeguarding privacy and minimizing ethical controversies.

### v2

@@ 4.1 Threat Model
Following previous work [21], we define a threat model to describe the goals and capabilities of the T2I service provider and the adversary, thereby clarifying the scope of our work.
Goals and Capabilities of the T2I Service Provider. The primary goal of the provider is to prevent unsafe generation and ensure the quality of the generated images in closed-source commercial scenarios. In terms of capabilities, the provider has complete access to the parameters of the T2I model, allowing them to optimize additional modules or update model parameters based on gradient information. Furthermore, the provider can integrate various safeguards, such as Negative Prompt, to improve the safety of the T2I model from a visual perspective.
Goals and Capabilities of the Adversary. The primary goal of the adversary is to bypass the safeguards in the target T2I model and induce the models to generate inappropriate high-quality content. We assume that the adversary has the capability to locally deploy an open-source T2I surrogate model, although this model fails to meet quality requirements. The adversary can utilize the local model to create or collect adversarial prompts, such as through optimization techniques. Additionally, we assume that the adversary interacts with the target model by submitting various prompt variants. However, the adversary only has black-box access and cannot directly access the internal parameters of the target model.
@@ 4.2 Overview
Based on the analysis of the properties of safety generation methods, our goal is to develop a safe generation framework that embodies these properties, providing comprehensive defense for T2I models in complex real-world scenarios to ensure that generated content adheres to usage policies.
Key Idea. T2I models are fundamentally text-driven image generators, with the core objective of producing high-quality images that are semantically consistent with the given prompt. Ensuring semantic alignment between the input text and the generated image is a fundamental requirement for T2I models. For well-trained and properly functioning models, generating unsafe content from anchor prompts—those without inappropriate descriptions—is highly challenging. For example, when prompted with “a photo of a cute dog,” the model is extremely unlikely to produce an image containing nudity, as this clearly violates semantic consistency. In other words, the presence of inappropriate content in generated images often originates from harmful descriptions in the input prompt, suggesting that the primary source of unsafe generation risks in T2I models lies in the text representation.
Based on this observation, we propose Embedding Sanitizer (ES), a vision-agnostic safe generation framework designed to erase inappropriate concepts directly from prompt embedding in a plug-and-play manner. This framework offers notable advantages in terms of compatibility. First, ES is decoupled from the vision modules and requires access only to the text encoder during training, allowing it to be flexibly integrated into any T2I model that shares the same text encoder, without requiring any modification to the underlying model. Second, while most existing safety approaches focus on removing inappropriate concepts from visual representations, usually by modifying the denoiser or denoising process, ES targets the removal of inappropriate concepts from the textual embedding itself. This makes it enabling seamless integration with various safeguards to mitigate unsafe generation at the source.
Next, we formalize the key idea of ES. Assume that the T2I model is equipped with a text encoder denoted as $\mathcal{F}_{t}$. Given a toxic prompt $P_{t}$ that contains inappropriate descriptions, such as ”A nude person”, ES aims to transform this into an anchor prompt $P_{a}$ with similar semantics, such as ”A dressed person”. This can be represented as:
$\text{ES}(\mathcal{F}_{t}(P_{t}))\approx\mathcal{F}_{t}(P_{a}),$ (1)
where ES takes the toxic prompt embedding $\mathcal{F}_{t}(P_{t})$ as input and sanitizes it to approximate the embedding of $P_{a}$.
@@
Overall Pipeline. The entire execution process of ES is illustrated in  Figure 2. At a high level, when a user inputs a prompt to query the T2I model, the text encoder first uses a tokenizer and embedding lookup to encode each token in the text prompt into embeddings. The text encoder then further processes these embeddings through a self-attention mechanism to extract high-level feature representations, which are fed as input to ES. Finally, ES performs a sanitization process on the embeddings, aiming to erase inappropriate concepts and transform them into anchor embeddings.
At a microscopic level, the embeddings output by the text encoder are first processed by EmbeddingNet (E-Net), which is designed to capture toxic embeddings containing inappropriate concepts within the original suspicious embeddings. Taking the example of ”a nude person,” the token ”nude” is the main element that leads to unsafe generation. E-Net outputs a highly toxic embedding (represented by the dark red box) for ”nude”. In contrast, for benign tokens such as ”a” and ”person,” E-Net outputs either non-toxic or mildly toxic embeddings (represented by the light red box). Next, ScoreNet (S-Net) takes the toxic embeddings as input and assigns a score to each token within the range [0, 1] based on the degree of toxicity. For toxic tokens, S-Net tends to predict higher scores to strongly suppress the target token. For benign tokens, S-Net tends to assign lower scores, aiming to minimize interference with benign concept generation. The scores predicted by S-Net are multiplied with the toxic embeddings predicted by E-Net, and suspicious concepts undergo sanitization by subtracting the toxic components from the suspicious embeddings in a residual connection. The sanitized toxic embeddings can then serve as an approximate representation of benign embeddings, providing guidance for generating appropriate content.
4.3 Model Architectural Design
As shown in Figure 2, the architecture of ES comprises two main components: E-Net and S-Net, which work together to perform sanitization functions. Below, we provide a detailed description of the architectural design of each component and explain how these designs confer interpretability and controllability to ES. The model hyperparameter configurations are summarized in Table II.
@@ 4.3 Model Architectural Design
E-Net. The design of E-Net is based on the Transformer encoder architecture and includes six attention modules to efficiently capture potential inappropriate content in the input prompt. Each attention module employs an 8-head multi-head attention mechanism to enable E-Net to semantically analyze various parts of the input, allowing for a deeper understanding of tokens containing inappropriate concepts. The role of E-Net is to capture harmful concepts from the original suspicious embedding, producing a set of toxic embeddings that represent inappropriate content. Suppose $P_{s}$ is a suspicious prompt that potentially contains inappropriate concepts and is encoded by the text encoder $\mathcal{F}_{t}$ to obtain the initial suspicious embedding, denoted as $\text{Emb}_{s}=\mathcal{F}_{t}(P_{s})$. E-Net, represented by $\mathcal{F}_{\theta}$, takes these embeddings as input and extracts identifiable toxic embeddings $\text{Emb}_{t}$, which can be formulated as:
@@
S-Net. S-Net is a multilayer perceptron (MLP) composed of two fully connected layers. Its input dimension matches the output dimension of the embedding generated by E-Net, and it applies a sigmoid activation function to constrain its output within [0, 1]. The primary function of S-Net is to assign a score to each token in the input prompt, evaluating its potential harm and indicating its contribution to the generation of inappropriate content, thus providing ES with interpretability. Ideally, S-Net assigns high scores to toxic tokens, such as nude, enabling stronger erasure of toxic embeddings during the subsequent sanitization step. On the other hand, it assigns low scores to benign tokens, such as dog, minimizing disruption to the generation of benign content. Represented by $\mathcal{F}_{\phi}$, S-Net takes the toxic embedding extracted by E-Net, $\text{Emb}_{t}$, as input and predicts a set of scores $\mathcal{S}$ for each token, expressed as:
@@
Embedding Sanitization. In the final sanitization step, the toxic embedding predicted by E-Net is multiplied by the predicted scores by S-Net, dynamically adjusting the sanitization strength according to each token’s score. This adjustment strategy provides two main advantages for ES ’s performance. First, it ensures that tokens identified as more toxic (i.e., with higher scores) undergo stronger sanitization, thereby enhancing the erasure of inappropriate content. Second, it preserves appropriate content by applying weaker sanitization to tokens identified as benign (i.e., with lower scores). Next, ES sanitizes the original suspicious concept by subtracting (erasing) toxic components from the suspicious embedding $\text{Emb}_{s}$ through residual connection to obtain an anchor embedding $\text{Emb}_{a}$ that no longer contains inappropriate concepts. The sanitization process is formulated as:
$\text{Emb}_{a}$ $=\mathcal{F}_{t}(P_{s})-\alpha\cdot\mathcal{F}_{\phi}(\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))\cdot\mathcal{F}_{\theta}(\mathcal{F}_{t}(P_{s}))$ (4)
@@
It is worth noting that we introduce a hyperparameter $\alpha$ to control the strength of sanitization in the inference phase. When $\alpha$ approaches 0, sanitization is weak, and ES ’s effect is minimal, resulting in $\text{Emb}_{a}$ being close to the original output of the text encoder. Conversely, as $\alpha$ increases, sanitization strength is significantly enhanced, effectively erasing harmful components of $\text{Emb}_{s}$ and producing a cleaner anchor embedding $\text{Emb}_{a}$. A smaller $\alpha$ reduces interference with the input prompt, preserving more of the original semantics but with weaker sanitization, while a larger $\alpha$ enhances sanitization at the cost of some semantic preservation. Thus, $\alpha$ provides ES with controllability, enabling it to adjust the sanitization strength flexibly according to the safety requirements of specific application scenarios, balancing between retaining content relevance and effectively erasing inappropriate content.
@@
4.4 Training Data Synthesis
@@
ES is essentially an end-to-end deep learning model that takes the original prompt embedding as input and outputs the sanitized anchor embedding, which in turn guides the T2I model to generate appropriate content. To sanitize prompt embeddings, we first select an anchor token corresponding to the target token to be erased, providing ES with a clear direction for sanitization. This approach enables us to frame the embedding sanitization as a transfer problem toward the anchor token. When the T2I model receives a prompt containing a target token, ES redirects this token to the predefined anchor token to achieve embedding sanitization. We design three approaches to construct $\langle$ target token, anchor token $\rangle$ pairs:
@@
$\bullet$ Antonyms of the target token, e.g., $\langle$ illegal, legal $\rangle$;
@@
$\bullet$ Hypernyms of the target token, e.g., $\langle$ blood, liquid $\rangle$;
$\bullet$ A neutral placeholder, e.g., $\langle$ weapon, __\__ $\rangle$.
@@
With the default settings, we use antonyms as anchor tokens to provide a more effective sanitization direction for the target token.
@@
Considering that the performance of deep learning models largely depends on vast amounts of training data, we design an automated method for synthesizing prompts to provide ES with an almost limitless supply of training samples. Specifically, we view the synthetic prompt $P$ as a sequence of $n$ tokens drawn from the vocabulary $\mathcal{V}$, represented as $P=\{\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n}\}$. During the data synthesis process, we first convert the target token set $\mathcal{C}_{t}=\{c_{t}^{1},c_{t}^{2},\ldots,c_{t}^{m}\}$ containing $m$ tokens and the anchor token set $\mathcal{C}_{a}=\{c_{a}^{1},c_{a}^{2},\ldots,c_{a}^{m}\}$ into their corresponding token IDs using a tokenizer. Next, within the maximum input length allowed by the text encoder (e.g., SD-v1.4 allows a maximum of 77 tokens), we randomly select a number of tokens to compose $P$, expressed as $P=\{(\text{token}_{1},\text{token}_{2},\ldots,\text{token}_{n})\mid\text{token}_{i}\in\mathcal{V},\forall i=1,2,\ldots,n,n\geq 3\}$. Here, the requirement that $n\geq 3$ ensures that $P$ contains at least three tokens: a starting token, a random token, and an ending token, which guarantees that $P$ complies with the input specifications and is not empty. Finally, we select random positions excluding the first and last tokens to replace them with random tokens from $\mathcal{C}_{t}$, thereby forming a toxic prompt containing a random number of target tokens, denoted as $P_{t}$. We then replace the target tokens in $P_{t}$ with their corresponding anchor tokens to create the anchor prompt $P_{a}$. This dataset construction strategy enables ES to learn how to transform toxic prompts $P_{t}$ into anchor prompts $P_{a}$, ensuring that the generated images reflect appropriate content without losing contextual relevance. The example in Figure 3 provides an intuitive illustration of a randomly synthesized $P_{t}$ and its corresponding $P_{a}$, where red tokens indicate target tokens and green tokens represent anchor tokens.
We adopt a bag-of-words prompt synthesis strategy for two main reasons. First, this approach offers high diversity and coverage across different combinations of target and anchor tokens, enabling the model to generalize across various contexts—including edge cases and adversarial expressions—and thereby enhancing ES’s robustness and adaptability in identifying inappropriate concepts. Second, existing T2I models are generally insensitive to the word order or syntactic structure of prompts [53], making the construction of semantically coherent natural language sentences unnecessary for effective embedding sanitization. Moreover, such coherent sentence generation introduces additional synthesis costs without providing clear benefits for ES. As validated in the ablation study in Section 7, the bag-of-words synthesis strategy provides stronger defense performance for ES compared to semantically coherent alternatives. Overall, automated training data synthesis significantly reduces the cost of manual annotation and mitigates the reliance on large-scale real-world datasets containing inappropriate content. This ensures that ES can be trained broadly while preserving privacy and minimizing ethical concerns.
4.5 Objective Function Formulation